OpenStudy (anonymous):

Let R be the relation xRy if and only if x^2+1=y^2+1. Since it is an equivalence relation, list all the elements of the equivalence classes [1], [0], [-1/4]

OpenStudy (anonymous):

I have already proven it's an equivalence relation but we didn't go over classes so much so I'm a bit confused

OpenStudy (anonymous):

\[x^2+1=y^2+1\iff x^2=y^2\iff x=\pm y\]

OpenStudy (anonymous):

therefore the equivalence class of \([1]\) is \(\{-1,1\}\)

OpenStudy (anonymous):

lot easier than the last one!

OpenStudy (anonymous):

oh ok! So for [0] the only element is 0?

OpenStudy (anonymous):

what about [-1/4]?...would it just be {-1/4, 1/4}?

myininaya (myininaya):

looks good

OpenStudy (anonymous):

as judge judy would say "the questions get harder"

OpenStudy (anonymous):

so my answers are correct then?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you@

OpenStudy (anonymous):

yw